OpenStudy (anonymous):

Need to calculate the Area under the "f(x)=e^-x²" between the points where y=0.5. I found out that the x value here is "±sqrt(ln(2))". Now I did the following: Integrate from -sqrt(ln(2)) to sqrt(ln(2)) f(x)=e^-x² However, If I look up the answer it says: Integrate from -sqrt(ln(2)) to sqrt(ln(2)) f(x)=e^(-x²)-(1/2) Why the -(1/2)?

OpenStudy (anonymous):

Mine \[\int\limits_{-\sqrt{\ln 2}}^{\sqrt{\ln 2}}e ^{-x²}\] The correct one: \[\int\limits_{-\sqrt{\ln 2}}^{\sqrt{\ln 2}}e ^{-x²}-0.5\]

OpenStudy (anonymous):

Okay you are finding the area of the region between y=e^-x^x and y=.5

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=Plot%5Be^%28-x%29^2%2C0.5%2C {x%2C-2%2C2}%5D

OpenStudy (anonymous):

Just doing integral of e^-x² find you area of region between the function and y=0 which is not what you want

OpenStudy (anonymous):

its - 1/2 because you are calculating the area inside or in between y=e^-x^2 and y=0.5

OpenStudy (anonymous):

Ah, that makes sense. Thanks guys!